OpenStudy (anonymous):

The area of this cylinder's base is one hundred fifty square inches. The height of the cylinder is 6 inches. What is the volume of the cone that has the same base area and height? Show the steps of your solution and explain your work.

OpenStudy (anonymous):

@carloshdzz23

OpenStudy (anonymous):

@Smokeybaby0417

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

\(Volume ~of ~cone = \pi r^2 \frac{h}{3}\) h = height r = radius

OpenStudy (anonymous):

@Phebe can you please help

OpenStudy (anonymous):

https://answers.yahoo.com/question/index?qid=20100919174612AAYLyRh this could help you